Find the number of nonzero solutions of the equation: z^2+2z̅=0.
Answer by stanbon(75887)

Find the number of nonzero solutions of the equation: z^2+2z̅=0.
------
If you are just looking for the number of solutions, use
the discriminant:
b^2-4ac = 2^2-4*1*0 = 4
---
Since it is > 0 there are 2 solutions.
====================
If you need the solutions:
Factor z^2+2z = 0
z(z+2) = 0
z = 0 or z = -2
====================
Cheers,
Stan H.